Ticket Number: HEA-NOC/20080208-11 Ticket Status: CLOSE
Ticket Type: unscheduled Resolver: GEANT
Ticket Opened: 20080208 16:45 UTC Problem Start: 20080208 16:05 UTC
Ticket Closed: 20080211 09:12 UTC Problem End: 20080208 17:39 UTC
Site/line: Géant STM-16 Secondary
At 16:05 the v4 & v6 backup BGP peerings to Geant dropped.
Traffic is still flowing as expected via the primary link (but with
asymmetric traffic as described in 20080205-13), but there is now no
resilience on these links.
20080208 16:42 BN Geant Service Desk say no work has been performed on
their side. No changes have been made on the HEAnet side. Investigating.
20080211 09:10 BN This outage was caused by a misconfiguration on the
Geant equipment. This was fixed at 17:39 on 20080208 and has remained
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20080208-11
HEAnet Limited. Registered in Ireland, No. 275301.